OpenStudy (anonymous):

really need help: A line has a rise of 3 and a run of 12. What is the slope?

OpenStudy (anonymous):

its very simple. slope = dy/dx = rise/run = 3/12 = ...

OpenStudy (anonymous):

yeah but like with numbers how could i find it out because this is so confusing.

OpenStudy (anonymous):

A line has a rise of 3 and a run of 12. What is the slope? 1/3 1/4 4

OpenStudy (johnweldon1993):

\[\large slope = \frac{rise}{run}\] you have rise = 3 and run = 12 up there so \[\large slope = \frac{3}{12}\] that can be simplified...

OpenStudy (anonymous):

to 1/4 and that is the answer...?

OpenStudy (anonymous):

yes, slope is 1/4

OpenStudy (anonymous):

tahnks guys cool